find the slope of the line that passes thru the given points:
(-8, 2) and (-3, 4)

Answer by KnightOwlTutor(293)

The standard linear equation is y=mx+b. In order to find the slope of the line(m) you must find the difference of the y values/difference of X values
Y2-y1/X2-X1 4-2/-3-(-8)=2/5 therefore the general equation is y=2/5x + b
In order to find out what b is you need the plug in the x,y values of a point.
ie (-8,2) 2=(2/5)(-8) + b b=26/5 to confirm this answer plug in the values of the second point and see if both sides equal.
4=(-3)(2)/5 + 26/5 -6/5 + 26/5= 20/5=4
the final equation is y=2/5x + 26/5.